# Variational Prior Replacement in Bayesian Inference and Inversion Xuebin Zhao1 and Andrew Curtis1

### v1

@@
Variational Prior Replacement in Bayesian Inference
and Inversion
@@ 1 School of GeoSciences, University of Edinburgh, United Kingdom
Bayesian Inference, Probability distributions, and Seismic TomographyMany scientific investigations require that the values of a set of model parameters are estimated using recorded data – a process referred to as inference. In Bayesian inference, information from both observed data and prior knowledge that existed independently, is combined to update model parameters probabilistically. The result is represented by the so-called posterior probability distribution function. This describes the probability (density) of each possible set of parameter values according to their consistency with the data and prior information. Prior information is often described by a prior probability distribution. This represents our belief about the range of values that the variables can take, and their relative probabilities when considered independently of the current recorded data. Situations arise in which we wish to change prior information: (i) the subjective nature of prior information (your view about the results of previous studies may differ from ours), (ii) cases in which we wish to test different states of prior information as independent hypothesis tests, and (iii) information from new, independent studies may emerge during the course of an investigation so prior information may evolve over time. Estimating the solution to any single inference problem is usually computationally costly, as it typically requires thousands or millions of model samples to be drawn and simulation of the observed data set that would have been recorded if each sample was true. Therefore, recalculating the Bayesian inference solution every time prior information changes can be extremely expensive.
We develop a mathematical formulation that allows prior information to be changed in a solution using variational methods, without performing Bayesian inference on each occasion. In this method, existing prior information is removed from a previously obtained posterior distribution and is replaced by new prior information. We therefore call the methodology variational prior replacement (VPR). We demonstrate VPR using a 2D seismic full waveform inversion example, where VPR provides almost identical posterior solutions compared to those obtained by solving independent inference problems using different prior distributions. The former can be completed within minutes even on a laptop whereas the latter requires days of computations using high-performance computing resources. We demonstrate the value of the method by comparing the posterior solutions obtained using three different types of prior information: Uniform, smoothing and geological prior distributions.
@@ 1 Introduction
In a wide variety of scientific and engineering fields, researchers seek to estimate unknown (latent) parameters using observed data by solving an inverse or inference problem. By approximating the physical system, it is usually possible to calculate a forward function which estimates the synthetic data that would have been observed if any particular set of model parameter values were true, and this function commonly has unique values. However, direct inversion of this function is difficult if not impossible, due to uncertainties in the finite number of measurable data, and to the nonlinearity of the forward function (Boyd & Vandenberghe, 2004). Therefore, in practice solutions to inverse problems are non-unique (Mosegaard & Tarantola, 1995; Mosegaard & Sambridge, 2002; Tarantola, 2005; Valentine & Sambridge, 2023).
@@
Bayesian inference solves fully nonlinear, non-unique inverse problems under a probabilistic framework by seeking to define the family of all plausible solutions within model parameter space. The solution to Bayesian inference is described by the so-called posterior probability distribution function (pdf - either a probability density function for continuous variables, or a set of probabilities for discrete variables), obtained by updating prior information about model parameters with new information from observed data. It provides a statistical description of how consistent is each solution with the data and prior information, and allows uncertainties in the inverse problem solution to be estimated (Tarantola, 2005; Arnold & Curtis, 2018).
@@
Global search or sampling based methods are often used to solve Bayesian inference problems, where samples with non-zero posterior probability values are retained, sometimes in proportion to their probabilities, to build an ensemble of model solutions. These solutions are used to estimate statistical properties of the posterior distribution that characterise the solution uncertainty. Monte Carlo is one of the most frequently used sampling methods, including Metropolis-Hastings Markov chain Monte Carlo (MH-McMC – Metropolis et al., 1953; Hastings, 1970; Press, 1968; Mosegaard & Tarantola, 1995; Sambridge & Mosegaard, 2002), trans-dimensional Monte Carlo (Green, 1995; Malinverno, 2002; Sambridge et al., 2006; Bodin & Sambridge, 2009; Galetti et al., 2017; Zhang et al., 2018), gradient-based Monte Carlo methods (Welling & Teh, 2011; Girolami & Calderhead, 2011; Fichtner et al., 2019; Gebraad et al., 2020; Zhao & Sen, 2021; Biswas & Sen, 2022; de Lima et al., 2023a; Berti et al., 2023) and informed-proposal Monte Carlo (Khoshkholgh et al., 2021, 2022). Global search methods that do not use Markov chains have been developed to solve Bayesian problems, either using optimisation to search for the most probable solution such as in simulated annealing (Kirkpatrick et al., 1983; Sen & Stoffa, 2013; Zhao et al., 2022b) and genetic algorithms (Stoffa & Sen, 1991; Sambridge & Drijkoningen, 1992), or algorithms that characterise the posterior distribution such as the neighbourhood algorithm (Sambridge, 1999a, b), prior sampling (Meier et al., 2007a, b; Käufl et al., 2016; Mosser et al., 2020; Bloem et al., 2024), exact sampling (Propp & Wilson, 1996; Walker & Curtis, 2014a) and direct estimation of posterior pdfs without sampling (Nawaz & Curtis, 2016). However, all such sampling methods present deficiencies when applied to complex or high dimensional inference problems such as slow convergence (Atchadé & Rosenthal, 2005; Andrieu & Thoms, 2008), which implies that a large number of model samples and their corresponding forward simulations are required.
@@
In geophysics, variational inference was first applied to estimate subsurface geological facies and petrophysical parameters using seismic data (Nawaz & Curtis, 2018, 2019; Nawaz et al., 2020), where a mean-field approximation (which ignores correlations between parameters) is employed to simplify the mathematical formulation of the variational problem (Bishop, 2006; Kucukelbir et al., 2017). Since then, more advanced variational methods have been developed for different geophysical problems, such as travel time tomography (Zhang & Curtis, 2020; Zhao et al., 2021; Levy et al., 2022), seismic migration (Siahkoohi et al., 2021, 2023), seismic amplitude inversion (Zidan et al., 2022), earthquake hypocentre inversion (Smith et al., 2022), slip distribution inversion (Sun et al., 2023), full waveform inversion (Zhang & Curtis, 2021b; Bates et al., 2022; Wang et al., 2023; Lomas et al., 2023; Izzatullah et al., 2023; Zhao & Curtis, 2024b; Yin et al., 2024a), and experimental design (Strutz & Curtis, 2024).
@@
Most studies mentioned above, whether using random sampling or variational methods, focus on performing Bayesian inference efficiently and accurately given a specific set of observed data and fixed prior knowledge. Over recent years, researchers made use of neural networks and other machine learning architectures to implement efficient Bayesian inference in which the posterior pdf can be obtained rapidly for any newly observed dataset (Devilee et al., 1999; Meier et al., 2007a, b; Shahraeeni & Curtis, 2011; Shahraeeni et al., 2012; de Wit et al., 2013; Käufl et al., 2014, 2016; Earp & Curtis, 2020; Earp et al., 2020; Zhang & Curtis, 2021a; Mosher et al., 2021; Wang et al., 2022; Hansen & Finlay, 2022; Alyaev & Elsheikh, 2022; Grana et al., 2022; Guan et al., 2023). However, almost no publications consider situations where we want to change (update) the prior information used in a previously performed inference process, or where we have multiple plausible prior hypotheses to be tested for the same observed data. In such cases one might have to perform the inference repeatedly with different priors which would become extremely expensive in many applications, even though the data used in each individual process do not change.
@@
Walker & Curtis (2014b) introduced a method called prior replacement, which allows prior information to be changed rapidly in Bayesian inference without repeating the full inference procedure for each individual prior. This is achieved by dividing the obtained posterior distribution by the current prior pdf in an attempt to remove its effect, and then multiplying by a new prior pdf to inject (update) the results with new prior information. The method has been demonstrated to be effective for varying prior information when estimating rock physics parameters (clay volume and sandstone matrix porosity) using seismic impedance data. Walker & Curtis (2014b) used (semi-)analytic methods to perform prior replacement, which requires the calculation of integrals of probability distributions over the entire parameter space, making it difficult to calculate for high dimensional problems. Moreover, the analytic calculation is only applicable under stringent conditions: first, the existing posterior distribution is represented by a mixture of Gaussian distributions. And second, the old and new prior distributions should be Uniform, Gaussian, or (possibly) other probability distributions whose integral over the parameter space and whose multiplication and division by Gaussian distributions are analytically tractable, such that the replacement of the new posterior distribution can be calculated analytically.
@@
In this paper, we develop a prior replacement methodology under the framework of variational inference, hence the name variational prior replacement (VPR). VPR addresses (relaxes) the issues mentioned above, making it applicable for high dimensional and complicated Bayesian inference problems. The rest of this paper is organised as follows. In Section 2, we review the Bayesian inference and prior replacement concept developed in Walker & Curtis (2014b). To perform prior replacement more efficiently, we introduce variational inference and derive the variational prior replacement (VPR) framework. In Section 3, we demonstrate the proposed method in a seismic full waveform inversion example, in which we compare the results obtained using VPR with those found using independent Bayesian inference for each prior pdf. We demonstrate the effectiveness of the method by testing three different prior distributions using the same observed data. Finally, we provide a brief discussion and draw conclusions from this study.
@@ 2.2 Prior Replacement
Consider a situation where we have different types of prior information about model parameter , defined by prior probability distributions , , and so on, and we wish to calculate the posterior distributions corresponding to these priors. Such an array of prior distributions might originate from the views of different experts, or might invoke different assumptions about the structures and properties that might pertain to the model, perhaps representing a range of different hypotheses to be tested (e.g., Bloem et al., 2024). A straightforward strategy is to apply Bayes’ rule to each prior distribution, and solve independent (prior specific) Bayesian inverse problem whenever prior information changes. However, such a prior specific approach is not practically feasible, since it is already expensive to perform a single inference process, especially for high dimensional problems with large datasets such as typically occurs in seismic tomography and full waveform inversion (FWI) problems.
Alternatively, suppose we have obtained a posterior distribution for data based on some old prior information. When new prior information becomes available, we might remove the effect of the old prior information from that posterior distribution, then inject the new prior information. Thus we would obtain the new posterior pdf without explicitly applying Bayes’ rule a second time.
@@
The first equality is simply Bayes’ rule. In the context of Bayesian inference, we are usually supplied with a fixed dataset. Hence, in both new and old situations, we assume that the observed data is the same. On the right hand side, and are two constants which are independent of model vector according to equation 2. We define for later convenience. Term essentially takes the role of changing (replacing) the old prior by the new prior pdf in Bayesian inference, and states how we inject new prior information. Denote
@@
Equation 5 has a form that allows us to evaluate the new posterior distribution from the old one by updating prior information after Bayesian inference, assuming that we know both and , and that we can evaluate the normalisation constant . Therefore, the inference itself needs to be solved only once, no matter how many different types of prior information we wish to apply. However, prior replacement is valid only under one condition: the new prior must have zero probability where the old prior has zero (or very small) probability values to avoid a numerically unstable situation of dividing by zero (Walker & Curtis, 2014b). Intuitively, the old prior should be broader than the new prior.
Equations 4 and 5 state two main operations involved in prior replacement. The former is a typical Bayesian problem in which the (old) posterior pdf is evaluated given the observed data and old prior information. The latter can be viewed as a quasi-Bayesian problem in the sense that the new posterior pdf is obtained by combining information from three probability distributions – similar to Bayes’ rule, but without the need to calculate the new likelihood from scratch. Calculation of these expressions can be computationally expensive, so in the following two sections we introduce efficient methods for these two operations, respectively.
@@ 2.3 Variational Inference
The Bayesian problem stated in equation 4 can be solved using either random sampling or variational inference methods. Markov chain Monte Carlo (McMC) is a typical sampling method that generates an ensemble of samples according to the posterior distribution (as the number of samples tends to infinity) (Mosegaard & Tarantola, 1995). However, McMC can be practically expensive since the required number of samples increases exponentially with the dimensionality of the model vector – a concept referred to as the curse of dimensionality (Curtis & Lomax, 2001).
Variational inference is an alternative to McMC that can be more efficient in certain situations. In variational inference, we define a family of probability distributions with fixed (predefined) complexity, within which we select one member that best approximates the unknown posterior distribution. Therefore, variational inference solves Bayesian problems using optimisation rather than random sampling. The optimal distribution can be found by minimising the discrepancy between the variational and posterior distributions.
@@
Since is a constant and independent of , minimising the KL is equivalent to maximising the first two terms on the right hand side of equation 7. In addition, since KL, these two terms together act as a lower bound of the logarithmic evidence, and they are usually defined to be the evidence lower bound of (ELBO):
Evaluating ELBO is easier than evaluating KL since it does not explicitly require the evidence term to be calculated, which is computationally intractable. Therefore, the variational problem is often solved by maximising the ELBO. The optimisation result is a probability distribution with
@@
In variational inference, there is a trade-off when choosing the variational family: it needs to be sufficiently expressive to provide an accurate approximation to the (potentially complex) posterior distribution, yet simple enough for efficient optimisation. Different choices of the family often result in different variational distributions, and also different methods.

### v2

@@
Variational Prior Replacement in Bayesian Inference and Inversion
@@ 1 School of GeoSciences, University of Edinburgh, United Kingdom
Bayesian Inference, Probability Distributions, Seismic Tomography, Inverse Theory, Waveform InversionMany scientific investigations require that the values of a set of model parameters are estimated using recorded data. In Bayesian inference, information from both observed data and prior knowledge is combined to update model parameters probabilistically by calculating the posterior probability distribution function. Prior information is often described by a prior probability distribution. Situations arise in which we wish to change prior information during the course of a scientific project. However, estimating the solution to any single Bayesian inference problem is often computationally costly, as it typically requires many model samples to be drawn, and the data set that would have been recorded if each sample was true must be simulated. Recalculating the Bayesian inference solution every time prior information changes can therefore be extremely expensive. We develop a mathematical formulation that allows the prior information that is embedded within a solution, to be changed using variational methods, without recalculating the original Bayesian inference. In this method, existing prior information is removed from a previously obtained posterior distribution and is replaced by new prior information. We therefore call the methodology variational prior replacement (VPR). We demonstrate VPR using a 2D seismic full waveform inversion example, in which VPR provides similar posterior solutions to those obtained by solving independent inference problems using different prior distributions. The former can be completed within minutes on a laptop computer, whereas the latter requires days of computations using high-performance computing resources. We demonstrate the value of the method by comparing the posterior solutions obtained using three different types of prior information: uniform, smoothing and geological prior distributions.
@@ 1 Introduction
In a wide variety of scientific and engineering applications, researchers seek to estimate unknown (latent) parameters using observed data by solving an inverse or inference problem. By approximating the physical system, it is usually possible to calculate a forward function which estimates the synthetic data that would have been observed if any particular set of model parameter values were true, and this function commonly has unique values. However, direct inversion of this function is difficult if not impossible, due to uncertainties in the finite number of measurable data, and to the nonlinearity of the forward function (Boyd & Vandenberghe, 2004). Typically in practise, solutions to such inverse problems are non-unique (Mosegaard & Tarantola, 1995; Mosegaard & Sambridge, 2002; Tarantola, 2005; Valentine & Sambridge, 2023).
@@
Bayesian inference solves fully nonlinear, non-unique inverse problems under a probabilistic framework by seeking to define the family of all plausible solutions within model parameter space. The solution to Bayesian inference is described by the so-called posterior probability distribution function (pdf - either a probability density function for continuous variables, or a set of probabilities for discrete variables), obtained by updating prior information about model parameters with new information from observed data. It provides a statistical description of how consistent is each solution with both the data and prior information, and allows uncertainties in the inverse problem solution to be estimated (Tarantola, 2005; Arnold & Curtis, 2018).
@@
Global search or sampling based methods are often used to solve Bayesian inference problems. Samples of parameter values with non-zero posterior probability values are retained, sometimes in proportion to their probabilities, to build an ensemble of model solutions. These solutions are used to estimate statistical properties of the posterior distribution that characterise the solution uncertainty. Monte Carlo is one of the most frequently used sampling methods, including Metropolis-Hastings Markov chain Monte Carlo (MH-McMC – Metropolis et al., 1953; Hastings, 1970; Press, 1968; Mosegaard & Tarantola, 1995; Sambridge & Mosegaard, 2002), trans-dimensional Monte Carlo (Green, 1995; Malinverno, 2002; Sambridge et al., 2006; Bodin & Sambridge, 2009; Galetti et al., 2017), gradient-based Monte Carlo methods (Welling & Teh, 2011; Girolami & Calderhead, 2011; Fichtner et al., 2019; Gebraad et al., 2020; Zhao & Sen, 2021; Biswas & Sen, 2022; de Lima et al., 2023a; Berti et al., 2023) and informed-proposal Monte Carlo (Khoshkholgh et al., 2021, 2022). Global search methods that do not use Markov chains have been developed to solve Bayesian problems, either using optimisation to search for the most probable solution such as in simulated annealing (Kirkpatrick et al., 1983; Sen & Stoffa, 2013; Zhao et al., 2022b) and genetic algorithms (Stoffa & Sen, 1991; Sambridge & Drijkoningen, 1992), or algorithms that characterise the posterior distribution such as the neighbourhood algorithm (Sambridge, 1999a, b), prior sampling (Meier et al., 2007a, b; Käufl et al., 2016; Mosser et al., 2020; Bloem et al., 2024), exact sampling (Propp & Wilson, 1996; Walker & Curtis, 2014a) and direct estimation of posterior pdfs without sampling (Nawaz & Curtis, 2016). However, all such sampling methods present deficiencies when applied to complex or high dimensional inference problems such as slow convergence (Atchadé & Rosenthal, 2005; Andrieu & Thoms, 2008), which implies that a large number of model samples and their corresponding forward simulations are required.
@@
In geophysics, variational inference was first applied to estimate subsurface geological facies and petrophysical parameters using seismic data (Nawaz & Curtis, 2018, 2019; Nawaz et al., 2020), where a mean-field approximation (which ignores correlations between parameters) is employed to simplify the mathematical formulation of the variational problem (Bishop, 2006; Kucukelbir et al., 2017). Since then, more advanced variational methods have been developed for different geophysical problems, such as travel time tomography (Zhang & Curtis, 2020a; Zhao et al., 2021; Levy et al., 2022), seismic migration (Siahkoohi et al., 2021, 2023), seismic amplitude inversion (Zidan et al., 2022), earthquake hypocentre inversion (Smith et al., 2022), slip distribution inversion (Sun et al., 2023), full waveform inversion (Zhang & Curtis, 2021b; Bates et al., 2022; Wang et al., 2023; Lomas et al., 2023; Izzatullah et al., 2023; Zhao & Curtis, 2024b; Yin et al., 2024a), and experimental design (Strutz & Curtis, 2024).
@@
Most studies mentioned above, whether using random sampling or variational methods, focus on performing Bayesian inference efficiently and accurately given a specific set of observed data and fixed prior knowledge. Over recent years, researchers made use of neural networks and other machine learning architectures to implement efficient Bayesian inference in which the posterior pdf can be obtained rapidly for any newly observed dataset (Devilee et al., 1999; Meier et al., 2007a, b; Shahraeeni & Curtis, 2011; Shahraeeni et al., 2012; de Wit et al., 2013; Käufl et al., 2014, 2016; Earp & Curtis, 2020; Earp et al., 2020; Zhang & Curtis, 2021a; Mosher et al., 2021; Wang et al., 2022; Hansen & Finlay, 2022; Alyaev & Elsheikh, 2022; Grana et al., 2022; Guan et al., 2023; Sun & Williamson, 2024). However, almost no publications consider situations where we want to change (update) the prior information used in a previously performed inference process, or where we have multiple plausible prior hypotheses to be tested for the same observed data. In such cases one might have to perform the inference repeatedly with different prior distributions; this would become extremely expensive in many applications, even though the data used in each individual case do not change.
@@
Walker & Curtis (2014b) introduced a method called prior replacement, which allows prior information to be changed rapidly in Bayesian inference without repeating the full inference procedure for each individual prior pdf (on occasion below, we may refer to this simply as the prior). This is achieved by dividing the obtained posterior distribution by the current prior pdf in an attempt to remove the effect of the latter, and then multiplying the result by a new prior pdf to inject (update) the results with new prior information. The method was demonstrated to be effective for varying prior information when estimating rock physics parameters (clay volume and sandstone matrix porosity) using seismic impedance data. Walker & Curtis (2014b) used (semi-)analytic methods to perform prior replacement, which requires the calculation of integrals of probability distributions over the entire parameter space, making it difficult to calculate for high dimensional problems. Moreover, the analytic calculation is only applicable under stringent conditions: first, the existing posterior distribution is represented by a mixture of Gaussian distributions. And second, the old and new prior distributions should be uniform, Gaussian, or (possibly) other probability distributions whose integral over the parameter space and whose multiplication and division by Gaussian distributions are analytically tractable, such that the replacement of the new posterior distribution can be calculated analytically.
In this paper, we develop a prior replacement methodology under the framework of variational inference, hence the name variational prior replacement (VPR). VPR addresses (relaxes) the issues mentioned above, making it applicable for high dimensional and complicated Bayesian inference problems. We test and demonstrate the method on a full waveform inversion (FWI) problem. Until the last few years there was no published fully nonlinear Bayesian solution to any FWI problem that approached a practical scale, due to the huge computational cost, and associated theoretical and algorithmic challenges (Gebraad et al., 2020; Zhang & Curtis, 2020b). While to-date a single study has extended the method to a three dimensional case (Zhang et al., 2023), the computational issues to make this approach mainstream remain. Advances that deliver even approximate results using greatly reduced computation are therefore significant.
@@
The rest of this paper is organised as follows. In Section 2, we review the Bayesian inference and prior replacement concept developed in Walker & Curtis (2014b). To perform prior replacement more efficiently, we introduce variational inference and derive the variational prior replacement (VPR) framework. In Section 3, we demonstrate the method using a seismic full waveform inversion example, in which we compare the results obtained using VPR with those found using independent Bayesian inference for each prior pdf. We demonstrate the effectiveness of the method by testing three different prior distributions using the same observed data. Finally, we provide a brief discussion and draw conclusions from this study.
@@ 2.2 Prior Replacement
Consider a situation where we have different sets of prior information (e.g., different hypotheses, or differing beliefs held by different people) about model parameter , defined by prior probability distributions , , and so on, and we wish to evaluate the implications of these various priors by calculating the corresponding posterior distributions. Such an array of prior distributions might originate from the views of different groups of experts, or might invoke different assumptions about the structures and properties that might pertain to the model, perhaps representing a range of different hypotheses to be tested and discriminated (e.g., Bloem et al., 2024). A straightforward strategy is to apply Bayes’ rule to each prior distribution, and solve independent (prior specific) Bayesian inverse problem whenever prior information changes. However, such a prior specific approach is not practically feasible, since it is already expensive to perform a single inference process, especially for high dimensional problems with large datasets such as typically occurs in seismic tomography and FWI problems.
Alternatively, suppose we have obtained a posterior distribution for data based on one specific type of prior information. When additional prior information becomes available or when different prior hypotheses exist and need to be discriminated, we might remove the effect of the existing prior information from the current posterior distribution, then inject different prior information. Thus we would obtain the desired posterior pdf without explicitly applying Bayes’ rule a second time. Below we denote prior and posterior pdfs that have been considered or calculated previously as old pdfs and those obtained by updating the prior distribution in this way as new ones; the words old and new in this context do not refer to situations where we update information because additional data have been collected, but rather to the order in which different prior distributions are combined with information in a fixed data set.
@@
The first equality is simply Bayes’ rule. In both new and old distributions, we assume that the observed data are the same. On the right hand side, and are two constants which are independent of model vector according to equation 2, so we define for later convenience. Term essentially takes the role of changing (replacing) the old prior by the new prior pdf in Bayesian inference, and states how we inject new prior information. Denote
@@
Equation 5 has a form that allows us to evaluate the new posterior distribution from the old one by updating (replacing) prior information after Bayesian inference, assuming that we know both and , and that we can evaluate the normalisation constant . Using this formulation, there is no need to perform new likelihood evaluations, which is normally the most computationally expensive step in solving an inverse problem, no matter how many different prior distributions we wish to inject. However, prior replacement is valid only under one condition: the new prior must have zero (or in practise, very small) probability where the old prior has zero probability values to avoid a numerically unstable situation of dividing by zero (Walker & Curtis, 2014b). Intuitively, the support of the new prior pdf must be a subset of that of the old one.
Equations 4 and 5 define the two main operations involved in prior replacement. The former is the solution to a typical Bayesian problem in which the (old) posterior pdf is evaluated given the observed data and existing prior information. The latter can be viewed as a quasi-Bayesian problem in the sense that Bayes rule applied to the new prior distribution is implicit within the formula, and the new posterior pdf is obtained by combining information from three probability distributions – of similar form to Bayes’ rule, but without the need to re-calculate the likelihood function from scratch. Calculation of these expressions can nevertheless be computationally expensive, so in the following two sections we introduce efficient methods for each of these operations, respectively.
@@ 2.3 Variational Inference
The Bayesian posterior distribution in equation 4 can be estimated using either random sampling or variational inference methods. Markov chain Monte Carlo (McMC) is a typical sampling method that generates an ensemble of samples distributed according to the posterior distribution as the number of samples tends to infinity (Mosegaard & Tarantola, 1995). However, McMC can be expensive in practise since the required number of samples increases exponentially with the dimensionality of model vector – a concept referred to as the curse of dimensionality (Curtis & Lomax, 2001).
Variational inference is an alternative to McMC that can be more efficient in certain situations. In variational inference, we define a family of probability distributions with fixed (predefined) complexity, within which we select a member that best approximates the unknown posterior distribution. Therefore, variational inference solves Bayesian problems using optimisation rather than random sampling. The optimal distribution can be found by minimising the discrepancy between the variational and posterior distributions.
@@
Since is a constant and independent of , minimising KL is equivalent to maximising the first two terms on the right hand side of equation 7. In addition, since KL, these two terms together act as a lower bound on the logarithmic evidence, and they are usually defined as the evidence lower bound (ELBO) of :
Evaluating ELBO is easier than evaluating KL since it does not explicitly require the evidence term to be calculated, which is often computationally intractable. The variational problem is therefore often solved by maximising the ELBO. The optimisation result is a probability distribution with
@@
In variational inference, there is a trade-off when choosing the variational family: it needs to be sufficiently expressive to provide an accurate approximation to the (potentially complex) posterior distribution, yet simple enough for efficient optimisation. Different choices of the family often result in different variational distributions, and also in different algorithms.